x^3-5=3 could you help me solve this i have to school my work
Answer by mananth(16949)

x^3-5=3
add -3 to both sides
x^3-5-3= 3-3
x^3-8=0
This is similar to
a^3-b^3 = (a-b)(a^2+ab+b^2)
=(x-2)(x^2+2x+4)